# Linying Lv

### v1

@@ 3.2.1 Information Content of Quantitative and Qualitative Information
This model incorporates the out-of-sample expanding estimation of $CAR_{[-1,+1]}$ using report text embeddings as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
@@
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. In column (3), the inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% to 14.9%, underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by quantitative analyst revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the three-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}{rev}$ model alone achieves an $R^{2}$ of 8.9%, while the combined input model improves this to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. This stability in coefficients suggests that the market-valued analyst information is distinct from, and not merely chasing back on, past returns or recent news events.
@@ 3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the role of analysts surrounding corporate earnings announcements (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information events reinforce each other in sequence. I, therefore, further study the information content of sell-side analyst reports around this most significant corporate event.
@@
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. A finding consistent with Huang et al. (2018) is that an overwhelming number of analyst reports are released immediately following the earnings announcements. Consistent with Huang et al. (2018), a large number of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks, and does not recover thereafter.
@@
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports during earnings announcement weeks. To account for potential temporal variation, I train Ridge models using separate samples for periods surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports relative to earnings announcements. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained outside these periods. Additionally, the $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen. This trend suggests that analysts react swiftly to earnings announcements, with the information content of their reports declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
@@
One potential concern in assessing the value of analyst reports is that analysts might merely be repeating the content of earning announcement calls, with the earnings announcement itself driving market reactions. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts data to differentiate the information content between the two text sources. The analysis focuses on reports released within a one-day window following earnings announcements, as this is when analysts are most likely to piggyback on the announcements.
@@
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcripts as input. [^11] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. with a Diebold-Mariano (DM) t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods of both texts. The input comprised the mean of all report text embeddings on the day following the earnings announcement. This firm-time level Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
@@
Incorporating both analyst report embeddings and earnings conference call embeddings in the Ridge regression resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
@@ 3.2.3 What Content in Analyst Reports is Valued by Market Investors?
Given the comprehensive discussion in equity reports, what specific content in analyst reports is the most valuable to the market? I employ the Shapley value approach in Section 2.4 to attribute the $R^{2}_{\mathrm{OOS}}$ of the full text to 17 topics. In this section, I use the token-weighted average of each sentence embedding as the report embedding. This sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content in analyst reports.
@@
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged into a single importance measure for each topic. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring a comprehensive attribution of explanatory power. Figure 2 ranks the Shapley values of 17 topics. Topics are ordered from left to right, with the most important topics on the left and the least important on the right. Each topic is represented by a bar, with a color gradient (from lightest to darkest) indicating the magnitude of its importance in terms of Shapley values.
@@
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each comprising 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
@@
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses, underlining the consistency of the results.
@@
Figure A5 in the Online Appendix reveals how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. The annual variations align with the full-context embedding results in Table 2, exhibiting a consistent story across different text embeddings. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite this volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable, indicating that analyst discussions of income statements consistently provide valuable information, even during market disruptions.
Figure A6 further supports this by showing a consistent pattern of key topics across industries. This suggests that the market values Income Statement Analyses and Financial Ratio Analyses universally, beyond specific sectors.
@@
To investigate whether this topic ranking is driven by analysts’ expertise in certain topics (supply effect) or investors’ differential responses to different topics (demand effect), Figure A7 compares Shapley values across analysts with varying experience levels and from different-sized brokerage firms. The remarkable consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) eliminates the concern that the topic ranking merely reflects analysts’ varying skill levels in different topics.
@@
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from the literature, particularly Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
@@
To facilitate this analysis, I further categorize sentences within the Income Statement Analyses topic into two sets of sub-topics: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a sophisticated prompt-based classification system, as detailed in Prompt 3:
@@
Figure 3 offers a detailed breakdown of the $R^{2}_{\mathrm{OOS}}$ across 18 topics, with the Income Statement category further divided into sub-topics. Specifically, the income interpretation subtopic contributes approximately three times more to the $R^{2}_{\mathrm{OOS}}$ than income acquisition, underscoring the market’s preference for context and explanation over mere data reporting. Additionally, the income realization subtopic also contributes about three times more to the $R^{2}_{\mathrm{OOS}}$ than income expectation, indicating a stronger market preference for analysis of actual financial results over forward-looking projections. These findings suggest that the market values analysts’ interpretation of realized income to a greater extent than their ability to simply report public information or make expectations about future performance.
@@ 3.2.4 Comparison of Revision Reports and Reiteration Reports
Table 5 summarizes the out-of-sample performance of numerical and textual inputs in both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. When recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information sets further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding aligns with the literature that attributes large stock price changes to analyst recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), also reveal a clear contrast between revisions and reiterations. While target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ of 20.88%, t-stat = 11.18).
Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. Interestingly, the $R^{2}_{\mathrm{OOS}}$ for report text in earnings forecast reiteration reports is negative, indicating no explanatory power for contemporaneous stock returns. However, when focusing on analyst earnings forecast revisions,, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination of these inputs yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.

### v2

@@ 3.2.1 Information Content of Quantitative and Qualitative Information
This model incorporates the out-of-sample estimation of $CAR_{[-1,+1]}$ using report text embeddings or forecast revisions as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
@@
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}_{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. The inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% in column (1) to 14.9% in column (3), underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by forecast revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the 3-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}_{rev}$ model alone achieves an $R^{2}$ of 8.9%, while combined input improves it to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. The stability of the coefficients suggests that market-valued analyst information is distinct from recent earnings news and is not simply a reaction to past returns.
@@ 3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the relationship between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information sources reinforce each other in sequence. I examine the information content of sell-side analyst reports around earnings announcements, the most significant corporate events.
@@
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. Consistent with Huang et al. (2018), a large portion of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks.
@@
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports surrounding earnings announcements. To account for potential temporal variation, I train Ridge models using separate samples of analyst reports released surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained beyond these periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen, declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week. This trend suggests that the market values information from analysts who react swiftly to earnings announcements.
@@
One potential concern is that analysts may simply restate information from earnings calls, and that market reactions are driven by the earnings announcements themselves rather than analyst reports. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts to differentiate the information content between the two text sources. I analyze reports released within one day after earnings announcements, when analysts are most likely to simply restate earnings information.
@@
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcript embeddings as input. [^10] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. The input comprised the mean of all report text embeddings on the day following the earnings announcement. The Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
@@
Incorporating both analyst report embeddings and earnings conference call embeddings in Ridge regressions resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
@@ 3.2.3 What Content in Analyst Reports is Valued by Market Investors?
After showing that information in report text is significant, a key question arises: what specific content in analyst reports is the most valuable to the market? To systematically evaluate topic importance, I employ the Shapley value approach to attribute the text-based $R^{2}_{\mathrm{OOS}}$ to the 17 topics described in Section 2.1.1. I use sentence-segmented embeddings as report representation. The sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content across sentences.
@@
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged over the possible different permutations into a single importance measure. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring an efficient attribution of explanatory power.
@@
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each contributing to 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
@@
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses. Figure A5 reveals consistent topic rankings across industries.
@@
Figure A6 in the Online Appendix shows how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite the volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable. Analyst discussions of income statements consistently provide valuable information, even during market disruptions.
@@
The relative importance of topics could stem from analysts’ expertise in certain areas (supply effect) or from investors’ varying responses to different topics (demand effect). Figure A7 compares Shapley values across analysts of different experience levels and brokerage firm sizes. The consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) suggests that topic rankings are not driven by differences in analysts’ topical expertise.
@@
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
@@
To facilitate the analysis, I categorize Income Statement sentences based on two criteria: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
@@
Figure 3 provides a detailed breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, with Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw data. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on analysis of actual results than on forward-looking projections. These patterns demonstrate that markets particularly value analysts’ interpretation of realized income, more than their reporting of public information or future expectations.
@@ 3.2.4 Comparison of Revision Reports and Reiteration Reports
Table 5 summarizes the out-of-sample performance of models with numerical and textual inputs using both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. When stock recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions are considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003). While reports with target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. For models with text from earnings forecast reiteration reports as input, the $R^{2}_{\mathrm{OOS}}$ is insignificant. When focusing on analyst earnings forecast revisions, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.